Without referring to the fraction chart, write the equivalent fraction for each of the numbers given below.

Write an equivalent fraction for:

a. 1/4 b. 1/2 c. 6/8 d.2/4

e.2/2 f. 2/8 g.8/8 h.4/8

i.4/4


1
Expert's answer
2021-06-15T14:56:58-0400

We can obtain equivalent fractions by multiplying and dividing by the same number :-

a) "\frac{1}{4}:\frac{1}{4}×\frac{2}{2}=\frac{2}{8}"

or "\frac{1}{4}×\frac{3}{3}=\frac{3}{12}"


b) "\frac{1}{2}:\frac{1}{2}×\frac{4}{4}=\frac{4}{8}"


"or\frac{4}{8}×\frac{2}{2}=\frac{8}{16}"


c) "\frac{6}{8}:\frac{6}{8}×\frac{3}{3}=\frac{18}{24}"


"or \frac{6}{8}×\frac{2}{2}=\frac{12}{16}"


d) "\frac{2}{4}:\frac{2}{4}×\frac{8}{8}=\frac{16}{32}"


or "\frac{2}{4}×\frac{2}{2}=\frac{4}{8}"


e) "\frac{2}{2}:\frac{2}{2}×\frac{3}{3}=\frac{6}{6}"


"or\frac{2}{2}×\frac{8}{8}=\frac{16}{16}"


f) "\frac{2}{8}:\frac{2}{8}×\frac{2}{2}=\frac{4}{16}"


"or \frac{2}{8}×\frac{3}{3}=\frac{6}{24}"


g)"\frac{8}{8}" : "\frac{8}{8}×\frac{2}{2}=\frac{16}{16}"

"or\frac{8}{8}×\frac{3}{3}=\frac{24}{24}"


h) "\frac{4}{8}: \frac{4}{8}×\frac{2}{2}=\frac{8}{16}"


"or \frac{4}{8}×\frac{3}{3}=\frac{12}{24}"


i) "\frac{4}{4}: \frac{4}{4}×\frac{2}{2}=\frac{8}{8}"


"or\frac{4}{4}×\frac{3}{3}=\frac{12}{12}"